# fold/.style=folder, grow’=0, s sep=0.1pt, arr/.style=edge=-latex

### v3

@@ Abstract
This paper presents a deep reinforcement learning approach for synthesizing unitaries into quantum circuits. Unitary synthesis aims to identify a quantum circuit that represents a given unitary while minimizing circuit depth, total gate count, a specific gate count, or a combination of these factors. While past research has focused predominantly on continuous gate sets, synthesizing unitaries from the parameter-free Clifford+T gate set remains a challenge. Although the time complexity of this task will inevitably remain exponential in the number of qubits for general unitaries, reducing the runtime for simple problem instances still poses a significant challenge. In this study, we apply the tree-search method Gumbel AlphaZero to solve the problem for a subset of exactly synthesizable Clifford+T unitaries. Our approach can synthesize unitaries for up to five qubits generated from the set of randomized quantum circuits with up to $60$ gates. Furthermore, our inference times are around $30$ seconds on a single GPU on average, surpassing state-of-the-art algorithms QuantumCircuitOpt and MIN-T-SYNTH for higher qubit numbers. Our work provides a competitive baseline for synthesis algorithms to be developed in the upcoming years.
@@
Unitary synthesis, Reinforcement learning, MCTS, AlphaZero
@@ I Introduction
In contrast to other approaches that are trained for one specific unitary [13, 22], our agent can synthesize a wide range of input unitaries without the need for retraining, highlight the effectiveness and generalizability of our Gumbel AlphaZero-based approach. To the best of the authors’ knowledge, we are the first to scale RL-based unitary synthesis past the one- or two-qubit setting with a generalizing agent.
@@ III-A Classical Unitary Synthesis
The first heuristic algoroithm for synthesizing a T-count optimal unitary was developed by Mosca and Mukhopadhyay [11]. Their algorithm, MIN-T-SYNTH, prunes the search space by exploiting the structure of the channel representation of the unitary, achieving a conjectured polynomial complexity relative to the T-count. Despite its advancements, the algorithm’s exponential complexity relative to the number of qubits remains a challenge.
@@
Finally, the work in Ref. [44] used decision diagram representation of a unitary matrix [45] to synthesize exactly representable unitaries with the Clifford+T gate set. Their algorithm was able to synthesize unitaries with up to 15 qubits and millions of T-gates.
@@ III-B Reinforcement Learning-Based Unitary Synthesis
For approximate synthesis, Moro et al. pioneered the concept of synthesizing quantum circuits from fixed unitaries as a sequential decision-making process [12]. Employing DQN [54] and PPO [55], they successfully synthesized general single-qubit Haar-unitaries under the HRC efficient universal set [56] and a rotational gate set. Chen et al. extended this approach to two-qubit unitaries using a combination of DQN and A* search, albeit with a search complexity that grows exponentially with the number of gates required [57].
In this paper, we explore the application of reinforcement learning to unitary synthesis, aiming to extend its scalability to circuits with more qubits and greater depth. We benchmark our method against MIN-T-SYNTH [11] and QuantumCircuitOpt [41] using structured unitaries such as controlled-Hadamard, Toffoli, and Peres gates.
@@ VI Experiments
In this section, we present the experimental results of our RL agent for synthesizing unitaries and describe our setup in Section VI-A. We evaluate the agent on two types of unitaries: random unitaries in Section VI-B and structured unitaries in Section VI-C. We compare the performance of our agent with the state-of-the-art synthesis algorithms MIN-T-SYNTH [11] and QuantumCircuitOpt  [41].
@@ VI-A Implementation and Setup
Our environment code was fully implemented in JAX, leveraging its JIT (just-in-time) compilation to address the inherent complexities of the training and inference process and enable seamless agent parallelization on a GPU. To ensure correctness, the calculation of the unitary using our framework for a given circuit were cross-validated with the comprehensive quantum computing software Qiskit [60]. For Gumbel AlphaZero, we utilize an implementation from the DeepMind mctx library [21] for all of our experiments, which is also implemented in JAX.
@@
a) Training: Our agent’s training is conducted over two million time steps, utilizing a replay buffer with the capacity to store 100,000 transitions. Here, we distribute the data collection across 256 concurrent environment workers and compute the training step gradients using a batch size of 1024. As a result, each transition sample has an update-to-data ratio of four. We sample the gate count of target circuits uniformly from the range $[3,\beta]$. The upper limit $\beta$ is set to $60,60,60,40$ for 2,3,4,5 qubit models, respectively. We reduce $\beta$ as the number of qubits increases to alleviate the exponential growth in the size of the observation and action spaces. The gate set of our experiments is $\mathcal{G}=\{H,T,T^{\dagger},S,S^{\dagger},Z,CX\}$.
@@ VI-B Results for Random Unitaries
For our evaluation of randomized unitaries, we sample 50 target unitaries from randomly generated circuits per data point. We do this using rejection sampling, where we sample random circuits until we find 50 distinct target unitaries with a specific number of T-gates or total number of gates. Specifically, we sample circuits with 0 to 20 T-gates and 3 to 60 total gates for 2,3 and 4 qubits. For 5 qubits, we use the same range of T-gates but limit ourselves to 0 to 40 total gates. For comparison, we evaluate the same circuits using PyZX [63], an optimization heuristic based on ZX-calculus [64], which can reduce the number of T-gates in a circuit. The results of PyZX optimization can be seen as heuristic upper bound on the T-count of a unitary if the corresponding circuit is known. We also compare the total number of gates after PyZX optimisation and our agent’s synthesis.
@@
(a) Comparison of average number of T-gates in the synthesised circuit by our RL agent and the optimised circuit by PyZX.
(b) Distribution of unitaries our RL agent successfully ($C_{HS}<0.1$) synthesised out of the 50 unitaries.
@@
Figure 4: Our RL agent’s performance on synthesizing 50 unitaries, generated by randomly sampling Clifford+T gate set, building a circuit and calculating the corresponding unitary. Here, the number of T-gates for each unitary is fixed and ranges from 0 to 20 for 2, 3 and 4 qubits and 0 to 15 for 5 qubits. (a) Our agent can synthesize 2 and 3 qubit unitaries with similar average T-gate count as PyZX. PyZX performs slightly better on average for 4 and 5 qubit unitaries. (b) We see a slight decline in performance for higher T-gate count for 2 and 3 qubit unitaries, and a significant drop for 4 and 5 qubit unitaries. The observed decline in performance can be attributed to the imbalance in the training set distribution. Specifically, there are fewer instances of circuits with larger T-gate counts compared to those with smaller counts.
@@
In Fig. 4 we plot the performance of our RL agent in synthesizing random unitaries for a fixed number of T-gates. Fig. 4(a) shows the average and median number of T-gates obtained by our RL agent over 50 random unitaries. Here, the line $x=y$ denotes the case where the number of T-gates in the input and the output circuits are the same. In Fig. 4(b), we plot the fraction of unitaries the agent successfully synthesized. The success is determined by calculating $C_{HS}$ (Eq. 1), for each of the 50 instances. When $C_{HS}>0.1$, evaluation is considered failed and $C_{HS}<0.1$ is considered success [^2] Since PyZX is not a synthesis tool there is no possibility of failure for such small qubit count. PyZX is always able to either reduce the number of T-gates in the circuit or at least keep it the same.. We empirically found the threshold of $0.1$ to work best for our experiments. Consequently, we verified whether the results with $C_{HS}<0.1$ match the target unitary, which we found to be the case in all the experiments. The error band shown in Fig. 4(a) around the RL agent curve is the fraction of unsuccessful runs out of the 50 unitaries considered. We see that the error band broadens for a higher number of T-gates, showing a decrease in performance. In Fig. 5, we show the performance for synthesizing unitaries for a fixed number of total gates.
@@
(a) Comparison of the average total gate count in the synthesized circuit by our RL agent and the optimized circuit by PyZX.
(b) Distribution of unitaries our RL agent successfully ($C_{HS}<0.1$) synthesised out of the 50 unitaries.
@@
Figure 5: Our RL agent’s performance on synthesizing 50 unitaries, generated by randomly sampling Clifford+T gate set, building a circuit and calculating the corresponding unitary. Here, the total number of gates for each unitary is fixed and ranges from 3 to 60 for 2, 3 and 4 qubits and 3 to 40 for 5 qubits. (a) Our agent can synthesize 2,3 and 4 qubit unitaries with better average total gate count than PyZX, although we see a drop in performance for higher total gate count. Average total number of gates is similar for the case of 5 qubit unitaries. (b) We see a slight decline in performance for 2 and 3 qubit unitaries, and a significant drop for 4 and 5 qubit unitaries for larger gate count.
@@
We find that our RL agent can synthesize all the target unitaries for small T-gate count and total gate count. The decrease in performance for T-gate count, can be attributed to the skewed distribution of T-gate numbers in the training set, meaning there are fewer circuits with larger T-gate counts given the current upper limit of 60 total gates. The performance decline for larger number of qubits can be attributed to the growing number of possibilities for gate placement on the qubit lines along with higher dimensional observation vectors (number of elements of the unitary matrix). For 3,4, and 5-qubit unitaries, our RL agent performs similarly to PyZX in terms of the number of T-gates in the optimized circuit. The reduction in total gate count using our RL agent was much better than PyZX since it has been observed [65] that PyZX T-gate optimization can increase the total number of gates. This highlights the benefits of our two-level optimization (both on the number of T-gates and total gates) procedure, which targets both the number of T-gates and the total gate count.
@@ VI-C Evaluation on Structured Unitaries
Table I: Benchmarking our method on structured unitaries against MIN-T-SYNTH [11] and QCOpt  [41]. We compare the number of T-gates and the run-time of synthesis. The success ratio (as defined in Fig. 3 and Section VI-B (b)), in the column for our RL agent, provides the number of successful runs (‘-’ indicates algorithm was unsuccessful in finding the synthesis either due to time limit or no convergence to the solution). The example decompositions of some unitaries are given in [10].
Qubits Unitary Our RL agent MIN-T-SYNTH QCOpt
T Time(s) T Time(s) T Time(s)
2 CP 3 10.69 3 0.007 3 6.15
CV 3 9.82 - - 3 4.16
CY 0 9.50 0 0.224 3 1.89
CZ 0 9.41 0 0.330 0 0.1
SWAP 0 9.27 0 0.353 0 0.01
W 2 $\bm{9.62}$ 0 3221.21
CH 2 $\bm{16.23}$ 2 147.12
CT - - - - - -
3 Toffoli 7 25.92 7 4.99 7 10800
@@ 3 Toffoli
We compare our agent’s performance with other synthesis algorithms and depict the results in Table I. The table includes the number of qubits in the input circuit, the target unitary, the results obtained by our agent, and the results of MIN-T-SYNTH and QuantumCircuitOpt (QCOpt for short). For MIN-T-SYNTH, we used the code provided by the authors of Ref. [11]. While the evaluation was successful for some of our chosen examples, manual code adjustments would have been required, leading to ‘-’ entries in the table. For QCOpt, the method exceeded the time limit of three hours without finding a solution in the basic Clifford+T gate set, which are also denoted by ‘-’ in the table.
Our method successfully synthesized all the unitaries, except for the controlled-T gate, which was not synthesized by any of the algorithms. To the best of our knowledge, it is unknown whether an exact synthesis is even possible, without the use of ancilla qubits as per Theorem 2 in Ref. [39]. As per Refs. [10, 11], we our RL agent achieves optimal of T-count for all target unitaries except $U_{1}$, which was synthesized with one extra T-gate compared to MIN-T-SYNTH.
Both QCOpt and our RL agent require an explicitly defined input gate set, which for $n$-qubit unitary would require $\mathcal{O}(n^{2})$ possible ways to place CNOT gates and $\mathcal{O}(n)$ one-qubit gates unless we already have prior knowledge about the decomposition of the unitary. The difference between QCOpt and our method is that the mixed-integer programming (MIP) method that QCOpt uses becomes highly inefficient in terms of run-time under such scaling. In comparison, our agent only has to account for these possibilities during training.
Comparing the run-times of our agent with the heuristic methods, we observed that, while both MIN-T-SYNTH and QCOpt are faster for simple unitaries, our agent takes approximately fifteen seconds for synthesis. However, as we analyze the increase in run-time with the number of qubits, we notice a significant increase in synthesis time for the other methods, whereas our method remains relatively constant. Therefore, the most advantageous aspect of our RL algorithm is the inference time as we scale to higher qubits.

### v4

@@ Abstract
This paper presents a deep reinforcement learning approach for synthesizing unitaries into quantum circuits. Unitary synthesis aims to identify a quantum circuit that represents a given unitary while minimizing circuit depth, total gate count, a specific gate count, or a combination of these factors. While past research has focused predominantly on continuous gate sets, synthesizing unitaries from the parameter-free Clifford+T gate set remains a challenge. Although the time complexity of this task will inevitably remain exponential in the number of qubits for general unitaries, reducing the runtime for simple problem instances still poses a significant challenge. In this study, we apply the tree-search method Gumbel AlphaZero to solve the problem for a subset of exactly synthesizable Clifford+T unitaries. Our method effectively synthesizes circuits for up to five qubits generated from randomized circuits with up to $60$ gates, outperforming existing tools like QuantumCircuitOpt and MIN-T-SYNTH in terms of synthesis time for larger qubit counts. Furthermore, it surpasses Synthetiq in successfully synthesizing random, exactly synthesizable unitaries. These results establish a strong baseline for future unitary synthesis algorithms.
@@
Unitary synthesis, Reinforcement learning,
MCTS, AlphaZero
@@ I Introduction
In contrast to other approaches that are trained for one specific unitary [13, 22], our agent can synthesize a wide range of input unitaries without the need for retraining, highlighting the effectiveness and generalizability of our Gumbel AlphaZero-based approach. To the best of the authors’ knowledge, we are the first to scale RL-based unitary synthesis past the one- or two-qubit setting with a generalizing agent.
@@ III-A Classical Unitary Synthesis
The first heuristic algorithm for synthesizing a T-count optimal unitary was developed by Mosca and Mukhopadhyay [11]. Their algorithm, MIN-T-SYNTH, prunes the search space by exploiting the structure of the channel representation of the unitary, achieving a conjectured polynomial complexity relative to the T-count. Despite its advancements, the algorithm’s exponential complexity relative to the number of qubits remains a challenge.
@@
Finally, Synthetiq [44] is a recently published tool used to synthesize circuits over arbitrary finite gate sets using Simulated annealing (SA). The approach is to iteratively modify a randomly built circuit, until a given criterion is met. The tool is versatile in doing exact synthesis using Clifford+T and other custom finite gate sets, approximate synthesis, synthesis for partially specified operators and also use composite gates such as the Toffoli gate in the gate set for synthesis.
@@ III-B Reinforcement Learning-Based Unitary Synthesis
For approximate synthesis, Moro et al. pioneered the concept of synthesizing quantum circuits from fixed unitaries as a sequential decision-making process [12]. Employing DQN [53] and PPO [54], they successfully synthesized general single-qubit Haar-unitaries under the HRC efficient universal set [55] and a rotational gate set. Chen et al. extended this approach to two-qubit unitaries using a combination of DQN and A* search, albeit with a search complexity that grows exponentially with the number of gates required [56]. Further stride towards approximately synthesizing unitaries for more than 2 qubits using RL was done in Ref. [57], although this was not the main aim of the work.
In this paper, we explore the application of reinforcement learning to unitary synthesis, aiming to extend its scalability to circuits with more qubits and greater depth.
@@ VI Experiments
In this section, we present the experimental results of our RL agent for synthesizing unitaries and describe our setup in Section VI-A. We evaluate the agent on two types of unitaries: random unitaries in Section VI-B and structured unitaries in Section VI-C. We benchmark our agent against state-of-the-art synthesis algorithms MIN-T-SYNTH [11], QuantumCircuitOpt [60] and Synthetiq [44].
@@ VI-A Implementation and Setup
Our environment code was fully implemented in JAX [61], leveraging its JIT (just-in-time) compilation to address the inherent complexities of the training and inference process and enable seamless agent parallelization on a GPU. To ensure correctness, the calculation of the unitary using our framework for a given circuit was cross-validated with the comprehensive quantum computing software Qiskit [62]. For Gumbel AlphaZero, we utilize an implementation from the DeepMind mctx library [21] for all of our experiments, which is also implemented in JAX.
@@
a) Training: Each qubit model (2-5 qubits) is trained for different durations as shown in Table I. The training utilizes a replay buffer with the capacity to store 100,000 transitions. Here, we distribute the data collection across 256 concurrent environment workers and compute the training step gradients using a batch size of 1024. As a result, each transition sample has an update-to-data ratio of four. We sample the gate count of target circuits uniformly from the range $[3,\beta]$ where $\beta$ depends on the number of qubits as shown in Table I. We reduce $\beta$ for 5 qubits to alleviate the exponential growth in the size of the observation and action spaces. The gate set of our experiments is $\mathcal{G}=\{H,T,T^{\dagger},S,S^{\dagger},Z,CX\}$. We highlight that, although the initial training takes a considerable amount of time, a trained model can be subsequently used multiple times for synthesis in a compilation pipeline for any exactly synthesizable unitary.
Table I: Table of training times and range of sampling gates for the model we use to synthesize unitaries
Qubits
Number of time steps
(in millions)
Training
time (days)
Range of
sampling gates
2 2 14 [3,60]
3 1.6 10.33 [3,60]
4 1 7.5 [3,60]
5 1 6.66 [3,40]
@@ VI-B Results for Random Unitaries
For our evaluation of randomized unitaries, we sample 50 target unitaries from randomly generated circuits per data point. We do this using rejection sampling, where we sample random circuits until we find 50 distinct target unitaries with a specific number of T-gates or total gates. Specifically, we sample circuits with 0 to 20 T-gates and 3 to 60 total gates for 2, 3, and 4 qubits. For 5 qubits, we use the same range of T-gates but limit ourselves to 0 to 40 total gates. For comparison, we evaluate the same circuits using PyZX [65], an optimization heuristic based on ZX-calculus [66], which can reduce the number of T-gates in a circuit, and Synthetiq [44] which is also a unitary synthesis tool. The results of PyZX optimization can be seen as heuristic upper bound on the T-count of a unitary if the corresponding circuit is known. We also compare the total number of gates for the three methods. The run-time of PyZX for the number of qubits we consider is less than a second, which is expected since it is an optimization heuristic. So, we do not compare its run-time with our RL agent.
@@
(a) Comparison of average number of T-gates in the synthesised circuit by our RL agent, optimised circuit by PyZX and Synthetiq.
(b) Distribution of unitaries our RL agent and Synthetiq successfully ($C_{HS}<0.1$ and time-out = 400 s) synthesised.
@@
Figure 4: Our RL agent’s performance on synthesizing 50 unitaries, generated by randomly sampling Clifford+T gate set, building a circuit and calculating the corresponding unitary. Here, the number of T-gates for each unitary is fixed and ranges from 0 to 20 for 2, 3 and 4 qubits and 0 to 15 for 5 qubits. (a) Our agent can synthesize 2 and 3 qubit unitaries with similar average T-gate count as PyZX whereas it outputs slightly more number of T-gates for 4 and 5 qubits. Synthetiq on the other hand performs worse, and after around 10 T-gates the number of circuits synthesized drops significantly pulling down the average, giving a misleading impression of better T-gate count. (b) Here, we compare the successful synthesis of a given unitary with a given number of T-gates, as done by our RL agent and Synthetiq. We see a slight decline in performance for higher T-gate count for 2 and 3 qubit unitaries, and a significant drop for 4 and 5 qubit unitaries. The observed decline in performance can be attributed to the imbalance in the training set distribution. Specifically, there are fewer instances of circuits with larger T-gate counts compared to those with smaller counts.
@@
In Fig. 4 we plot the performance of our RL agent in synthesizing random unitaries for a fixed number of T-gates. Fig. 4(a) shows the average and median number of T-gates obtained by our RL agent over 50 random unitaries. Here, the line $x=y$ denotes the case where the number of T-gates in the input and the output circuits are the same. In Fig. 4(b), we plot the fraction of unitaries successfully synthesized by our agent and Synthetiq. The success for our RL agent is determined by calculating $C_{HS}$ (Eq. 1), for each of the 50 instances. When $C_{HS}>0.1$, evaluation is considered failed and $C_{HS}<0.1$ is considered success [^2] Since PyZX is not a synthesis tool there is no possibility of failure for such small qubit count. PyZX is always able to either reduce the number of T-gates in the circuit or at least keep it the same.. We empirically found the threshold of $0.1$ to work best for our experiments. Consequently, we verified whether the results with $C_{HS}<0.1$ match the target unitary, which we found to be the case in all the experiments. Success rate for Synthetiq is determined by a time-out condition, in our case, 400 seconds on a 48 core processor. If no circuit is found within this time-limit the synthesis is deemed a failure. The error band shown in Fig. 4(a) around the RL agent and Sythetiq curve is the fraction of unsuccessful runs out of the 50 unitaries considered. We see that the error band broadens for a higher number of T-gates, showing a decrease in performance. In Fig. 5, we show the performance for synthesizing unitaries for a fixed number of total gates.
@@
(a) Comparison of the average total gate count in the synthesized circuit by our RL agent, optimized circuit by PyZX and Synthetiq.
(b) Distribution of unitaries our RL agent and Synthetiq successfully ($C_{HS}<0.1$ and time-out = 400 s) synthesised.
@@
Figure 5: Our RL agent’s performance on synthesizing 50 unitaries, generated by randomly sampling Clifford+T gate set, building a circuit and calculating the corresponding unitary. Here, the total number of gates for each unitary is fixed and ranges from 3 to 60 for 2, 3 and 4 qubits and 3 to 40 for 5 qubits. (a) Our agent can synthesize 2, 3 and 4 qubit unitaries with better average total gate count than PyZX and Synthetiq, although we see a drop in performance for higher total gate count compared to PyZX. Average total number of gates is similar for the case of 5 qubit unitaries. (b) Here, we compare the successful synthesis of a given unitary with a given number of T-gates, as done by our RL agent and Synthetiq. We see a slight decline in performance for 2 and 3 qubit unitaries, and a significant drop for 4 and 5 qubit unitaries for larger gate count for our RL agent, whereas Synthetiq performs worse than RL for larger number of gates.
@@
We find that our RL agent can synthesize all the target unitaries for small T-gate count and total gate count. The decrease in performance for T-gate count, can be attributed to the skewed distribution of T-gate numbers in the training set, meaning there are fewer circuits with larger T-gate counts given the current upper limit of 60 total gates. The performance decline for larger number of qubits can be attributed to the growing combinatorics for gate placement on the qubit lines along with higher dimensional observation vectors (number of elements of the unitary matrix). For 3, 4, and 5-qubit unitaries, our RL agent performs similarly to PyZX in terms of the number of T-gates in the optimized circuit. The reduction in total gate count using our RL agent was much better than PyZX since it has been observed [67] that PyZX T-gate optimization can increase the total number of gates. This highlights the benefits of our two-level optimization (both on the number of T-gates and total gates) procedure, which targets both the number of T-gates and the total gate count. We also see the performance of Synthetiq against our RL agent on these randomly generated unitaries. The fraction of unitaries Synthetiq is able to synthesize degrades faster for deeper circuits compared to our RL agent both on the dataset of T-gates and total number of gates. For larger qubits and larger than 10 T-gates in the original circuit, the number of circuits Synthetiq is able to synthesise drops significantly as seen in Fig. 4(b) which pulls down the average (Fig. 4(a)), giving a misleading impression of better T-gate count in the synthesized circuit.
@@ VI-C Evaluation on Structured Unitaries
Table II: Benchmarking our method on structured unitaries against MIN-T-SYNTH [11], QCOpt  [60] and Synthetiq [44]. We compare the number of T-gates and the run-time of synthesis (‘-’ indicates algorithm was unsuccessful in finding the synthesis either due to time limit or no convergence to the solution). The example decompositions of some unitaries are given in [10].
Qubits Unitary Our RL agent MIN-T-SYNTH QCOpt Synthetiq
T Time(s) T Time(s) T Time(s) T Time(s)
2 CV 3 9.82 - - 3 4.16 3 0.089
CY 0 9.50 0 0.224 0 16.84 0 0.084
CZ 0 9.41 0 0.330 0 0.1 0 0.08
SWAP 0 9.27 0 0.353 0 0.01 0 0.08
W 2 9.62 - - 0 3221.21 2 0.087
CP 3 10.69 3 0.007 3 6.15 3 0.082
CH 2 16.23 - - 2 147.12 2 0.086
CT - - - - - - - -
3 Toffoli 7 25.92 7 4.99 7 10800 7 0.238
@@ 2 Toffoli
We compare our agent’s performance with other synthesis algorithms and depict the results in Table II. The table includes the number of qubits in the input circuit, the target unitary, the results obtained by our agent, and the results of MIN-T-SYNTH, QuantumCircuitOpt (QCOpt for short) and Synthetiq. For MIN-T-SYNTH, we used the code provided by the authors of Ref. [11]. While the evaluation was successful for some of our chosen examples, manual code adjustments would have been required, leading to ‘-’ entries in the table. For QCOpt, the method exceeded the time limit of three hours without finding a solution in the basic Clifford+T gate set, which are also denoted by ‘-’ in the table. For Synthetiq, we set a time limit of 800 seconds on a 48 core processor machine using all the cores, with ‘-’ in the table if the time-limit is exceeded.
Our method successfully synthesized all the unitaries, except for the controlled-T gate, which was not synthesized by any of the algorithms. From Corollary 2 in Ref. [26], we know that controlled-T cannot be synthesized exactly without the use of an ancilla qubit. As per Refs. [10, 11], our RL agent achieves optimal T-count for all target unitaries except $U_{1}$, which was synthesized with one extra T-gate compared to MIN-T-SYNTH. We also see that Synthetiq out-performs our method in all the instances except for the 5-qubit x-Mod 5, 1-bit adder and $U_{1}$ in the time for synthesis.
Both QCOpt and our RL agent require an explicitly defined input gate set, which for $n$-qubit unitary would require $\mathcal{O}(n^{2})$ possible ways to place CNOT gates and $\mathcal{O}(n)$ one-qubit gates unless we already have prior knowledge about the decomposition of the unitary. The difference between QCOpt and our method is that the mixed-integer programming (MIP) underlying QCOpt becomes highly inefficient in terms of run-time under such scaling. In comparison, our agent only has to account for these possibilities during training.
Comparing the run-times of our agent with the heuristic methods, we observed that, while MIN-T-SYNTH, QCOpt and Synthetiq are faster for simple unitaries, our agent takes approximately fifteen seconds for synthesis. However, as we analyze the increase in run-time with the number of qubits, we notice a significant increase in synthesis time for the other methods, whereas our method remains relatively constant. With Synthetiq, we see an increase in run-time by almost a factor of 10 with the number of qubits. Therefore, the most advantageous aspect of our RL algorithm is the inference time as we scale up till five qubits.